I am having major problems solving by completing the square 2x^2-6x=5. I have no idea where to begin
Answer by fractalier(6550)

The first step is to divide everything by 2...so we have
2x^2 - 6x = 5
x^2 - 3x = 5/2
Now continue with the standard procedure (except that this one involves yucky fractions)...lol...
x^2 - 3x + 9/4 = 5/2 + 9/4 (half of 3 squared is 9/4)
(x - 3/2)^2 = 10/4 + 9/4
(x - 3/2)^2 = 19/4 now take the square root of both sides
x - 3/2 = ± sqrt(19/4)
x - 3/2 = ± sqrt(19) / 2
x = (3 ± sqrt(19)) / 2